# Toward a Well-Calibrated Discrimination via Survival Outcome-Aware Contrastive Learning

### v1

@@ 4.3 Network Description
where $\beta$ is a balancing coefficient that trade-offs between the two loss terms; the effect of $\beta$ is provided in C.2. Please find the pseudo-code of ConSurv in H.
@@ Appendix B Augmentation Method
METABRIC NWTCO GBSG FLCHAIN
Methods CI IBS DDC D-CAL CI IBS DDC D-CAL CI IBS DDC D-CAL CI IBS DDC D-CAL
NI 0.650 0.096 0.052 0.003 0.704 0.059 0.250 0.137 0.681 0.157 0.070 0.000 0.788 0.088 0.247 0.096
SCARF 0.688 0.094 0.097 0.089 0.729 0.078 0.222 0.360 0.696 0.138 0.180 0.004 0.810 0.084 0.291 0.204
@@ C.1 Performance on Time-dependent C-index and Brier score
Table 5: Comparison of time-dependent C-index.
METABRIC NWTCO GBSG
Methods 25% 50% 75% 25% 50% 75% 25% 50% 75%
CoxPH $0.653_{\pm 0.024}$ $0.645_{\pm 0.018}$ $0.637_{\pm 0.018}$ $0.718_{\pm 0.026}$ $0.711_{\pm 0.025}$ $0.706_{\pm 0.026}$ $0.714_{\pm 0.027}$ $0.678_{\pm 0.019}$ $0.664_{\pm 0.018}$
DeepSurv $0.643_{\pm 0.034}$ $0.620_{\pm 0.024}$ $0.607_{\pm 0.029}$ $0.640_{\pm 0.081}$ $0.636_{\pm 0.079}$ $0.634_{\pm 0.078}$ $0.700_{\pm 0.056}$ $0.668_{\pm 0.047}$ $0.654_{\pm 0.043}$
DeepHit $0.605_{\pm 0.045}$ $0.659_{\pm 0.044}$ $0.603_{\pm 0.032}$ $0.717_{\pm 0.023}$ $\textbf{0.723}_{\pm\textbf{0.023}}$ $0.704_{\pm 0.024}$ $0.630_{\pm 0.045}$ $0.626_{\pm 0.030}$ $0.609_{\pm 0.028}$
DRSA $0.639_{\pm 0.030}$ $0.605_{\pm 0.020}$ $0.592_{\pm 0.022}$ $0.713_{\pm 0.020}$ $0.706_{\pm 0.019}$ $0.704_{\pm 0.019}$ $0.725_{\pm 0.687}$ $0.687_{\pm 0.015}$ $0.671_{\pm 0.016}$
DCS $0.646_{\pm 0.030}$ $0.615_{\pm 0.023}$ $0.606_{\pm 0.024}$ $0.675_{\pm 0.026}$ $0.657_{\pm 0.020}$ $0.643_{\pm 0.029}$ $0.720_{\pm 0.025}$ $0.687_{\pm 0.015}$ $0.674_{\pm 0.017}$
X-CAL $0.661_{\pm 0.025}$ $0.632_{\pm 0.018}$ $0.619_{\pm 0.019}$ $0.687_{\pm 0.026}$ $0.650_{\pm 0.029}$ $0.628_{\pm 0.034}$ $\textbf{0.727}_{\pm\textbf{0.025}}$ $0.680_{\pm 0.016}$ $0.676_{\pm 0.017}$
ConSurv $\textbf{0.686}_{\pm\textbf{0.028}}$ $\textbf{0.660}_{\pm\textbf{0.021}}$ $\textbf{0.645}_{\pm\textbf{0.023}}$ $\textbf{0.718}_{\pm\textbf{0.025}}$ $0.709_{\pm 0.024}$ $\textbf{0.705}_{\pm\textbf{0.025}}$ $0.723_{\pm 0.030}$ $\textbf{0.689}_{\pm\textbf{0.019}}$ $\textbf{0.676}_{\pm\textbf{0.020}}$
FLCHAIN SUPPORT SEER
Methods 25% 50% 75% 25% 50% 75% 25% 50% 75%
CoxPH $0.796_{\pm 0.019}$ $\textbf{0.793}_{\pm\textbf{0.015}}$ $0.790_{\pm 0.014}$ $0.603_{\pm 0.006}$ $0.607_{\pm 0.006}$ $0.609_{\pm 0.006}$ $0.865_{\pm 0.018}$ $0.890_{\pm 0.023}$ $0.855_{\pm 0.022}$
DeepSurv $0.787_{\pm 0.028}$ $0.783_{\pm 0.027}$ $0.780_{\pm 0.028}$ $0.603_{\pm 0.009}$ $0.605_{\pm 0.009}$ $0.608_{\pm 0.010}$ $0.838_{\pm 0.007}$ $0.835_{\pm 0.021}$ $0.836_{\pm 0.002}$
DeepHit $0.796_{\pm 0.020}$ $\textbf{0.793}_{\pm\textbf{0.014}}$ $0.785_{\pm 0.014}$ $0.516_{\pm 0.007}$ $0.489_{\pm 0.008}$ $0.521_{\pm 0.007}$ $0.847_{\pm 0.016}$ $0.841_{\pm 0.020}$ $0.827_{\pm 0.020}$
DRSA $0.776_{\pm 0.023}$ $0.776_{\pm 0.016}$ $0.772_{\pm 0.013}$ $0.586_{\pm 0.009}$ $0.583_{\pm 0.007}$ $0.580_{\pm 0.009}$ $0.840_{\pm 0.073}$ $0.839_{\pm 0.077}$ $0.820_{\pm 0.083}$
DCS $0.788_{\pm 0.024}$ $0.783_{\pm 0.019}$ $0.779_{\pm 0.018}$ $0.604_{\pm 0.008}$ $0.602_{\pm 0.009}$ $0.602_{\pm 0.009}$ $0.864_{\pm 0.019}$ $0.863_{\pm 0.023}$ $0.858_{\pm 0.023}$
X-CAL $0.793_{\pm 0.021}$ $0.788_{\pm 0.015}$ $0.784_{\pm 0.016}$ $0.610_{\pm 0.006}$ $0.608_{\pm 0.006}$ $0.608_{\pm 0.007}$ $0.840_{\pm 0.023}$ $0.839_{\pm 0.019}$ $0.833_{\pm 0.018}$
ConSurv $\textbf{0.796}_{\pm\textbf{0.022}}$ $0.791_{\pm 0.017}$ $\textbf{0.786}_{\pm\textbf{0.015}}$ $\textbf{0.616}_{\pm\textbf{0.006}}$ $\textbf{0.617}_{\pm\textbf{0.006}}$ $\textbf{0.618}_{\pm\textbf{0.006}}$ $\textbf{0.870}_{\pm\textbf{0.017}}$ $\textbf{0.866}_{\pm\textbf{0.020}}$ $\textbf{0.862}_{\pm\textbf{0.020}}$
@@
METABRIC NWTCO GBSG
Methods 25% 50% 75% 25% 50% 75% 25% 50% 75%
CoxPH $0.167_{\pm 0.009}$ $\textbf{0.215}_{\pm\textbf{0.014}}$ $\textbf{0.236}_{\pm\textbf{0.038}}$ $\textbf{0.111}_{\pm\textbf{0.009}}$ $0.114_{\pm 0.008}$ $0.114_{\pm 0.008}$ $\textbf{0.145}_{\pm\textbf{0.010}}$ $0.211_{\pm 0.007}$ $0.219_{\pm 0.010}$
DeepSurv $0.173_{\pm 0.010}$ $0.228_{\pm 0.016}$ $0.249_{\pm 0.048}$ $0.121_{\pm 0.011}$ $0.124_{\pm 0.011}$ $0.125_{\pm 0.012}$ $0.146_{\pm 0.011}$ $0.212_{\pm 0.009}$ $0.221_{\pm 0.013}$
DeepHit $0.181_{\pm 0.009}$ $0.256_{\pm 0.010}$ $0.263_{\pm 0.026}$ $0.151_{\pm 0.009}$ $0.143_{\pm 0.008}$ $0.154_{\pm 0.009}$ $0.168_{\pm 0.012}$ $0.242_{\pm 0.008}$ $0.250_{\pm 0.006}$
DRSA $0.264_{\pm 0.041}$ $0.343_{\pm 0.038}$ $0.337_{\pm 0.079}$ $0.281_{\pm 0.030}$ $0.294_{\pm 0.039}$ $0.310_{\pm 0.051}$ $0.210_{\pm 0.016}$ $0.322_{\pm 0.022}$ $0.349_{\pm 0.020}$
DCS $0.190_{\pm 0.020}$ $0.265_{\pm 0.038}$ $0.293_{\pm 0.083}$ $0.119_{\pm 0.010}$ $0.107_{\pm 0.011}$ $0.122_{\pm 0.024}$ $0.153_{\pm 0.010}$ $0.210_{\pm 0.006}$ $0.218_{\pm 0.010}$
X-CAL $0.180_{\pm 0.012}$ $0.234_{\pm 0.021}$ $0.254_{\pm 0.057}$ $0.113_{\pm 0.009}$ $\textbf{0.106}_{\pm\textbf{0.013}}$ $0.137_{\pm 0.036}$ $0.150_{\pm 0.011}$ $\textbf{0.207}_{\pm\textbf{0.007}}$ $\textbf{0.217}_{\pm\textbf{0.012}}$
ConSurv $\textbf{0.166}_{\pm\textbf{0.009}}$ $0.221_{\pm 0.015}$ $0.255_{\pm 0.045}$ $0.111_{\pm 0.009}$ $0.113_{\pm 0.008}$ $\textbf{0.113}_{\pm\textbf{0.009}}$ $0.146_{\pm 0.009}$ $0.209_{\pm 0.008}$ $\textbf{0.217}_{\pm\textbf{0.010}}$
FLCHAIN SUPPORT SEER
Methods 25% 50% 75% 25% 50% 75% 25% 50% 75%
CoxPH $\textbf{0.074}_{\pm\textbf{0.006}}$ $\textbf{0.112}_{\pm\textbf{0.006}}$ $\textbf{0.133}_{\pm\textbf{0.007}}$ $\textbf{0.219}_{\pm\textbf{0.003}}$ $0.197_{\pm 0.008}$ $0.180_{\pm 0.009}$ $\textbf{0.005}_{\pm\textbf{0.001}}$ $\textbf{0.009}_{\pm\textbf{0.001}}$ $\textbf{0.014}_{\pm\textbf{0.001}}$
DeepSurv $0.076_{\pm 0.007}$ $0.114_{\pm 0.009}$ $0.137_{\pm 0.011}$ $0.220_{\pm 0.004}$ $0.198_{\pm 0.009}$ $0.180_{\pm 0.010}$ $\textbf{0.005}_{\pm\textbf{0.001}}$ $0.010_{\pm 0.001}$ $0.016_{\pm 0.001}$
DeepHit $0.096_{\pm 0.005}$ $0.181_{\pm 0.004}$ $0.244_{\pm 0.004}$ $0.340_{\pm 0.005}$ $0.282_{\pm 0.003}$ $0.240_{\pm 0.005}$ $0.016_{\pm 0.000}$ $0.020_{\pm 0.001}$ $0.026_{\pm 0.001}$
DRSA $0.141_{\pm 0.013}$ $0.251_{\pm 0.027}$ $0.323_{\pm 0.031}$ $0.285_{\pm 0.013}$ $0.269_{\pm 0.022}$ $0.249_{\pm 0.021}$ $0.011_{\pm 0.006}$ $0.021_{\pm 0.016}$ $0.034_{\pm 0.026}$
DCS $0.079_{\pm 0.008}$ $0.119_{\pm 0.008}$ $0.147_{\pm 0.014}$ $0.222_{\pm 0.006}$ $0.209_{\pm 0.015}$ $0.202_{\pm 0.017}$ $\textbf{0.005}_{\pm\textbf{0.001}}$ $0.010_{\pm 0.001}$ $0.015_{\pm 0.002}$
X-CAL $0.080_{\pm 0.005}$ $0.117_{\pm 0.007}$ $0.144_{\pm 0.012}$ $0.220_{\pm 0.005}$ $0.207_{\pm 0.016}$ $0.201_{\pm 0.018}$ $0.006_{\pm 0.001}$ $0.011_{\pm 0.002}$ $0.019_{\pm 0.003}$
ConSurv $0.077_{\pm 0.006}$ $0.115_{\pm 0.007}$ $0.136_{\pm 0.008}$ $\textbf{0.219}_{\pm\textbf{0.005}}$ $\textbf{0.195}_{\pm\textbf{0.010}}$ $\textbf{0.177}_{\pm\textbf{0.010}}$ $\textbf{0.005}_{\pm\textbf{0.001}}$ $\textbf{0.009}_{\pm\textbf{0.001}}$ $0.015_{\pm 0.001}$
C.2 Sensitive Analysis

### v2

@@ 4.3 Network Description
where $\beta$ is a balancing coefficient that trade-offs between the two loss terms; the effect of $\beta$ is provided in C.3. Please find the pseudo-code of ConSurv in H.
@@ Appendix B Augmentation Method
METABRIC NWTCO
Method CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$ CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$
NI 0.650 0.096 0.052 0.003 0.704 0.059 0.250 0.137
SCARF 0.688 0.094 0.097 0.089 0.729 0.078 0.222 0.360
GBSG FLCHAIN
Method CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$ CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$
NI 0.681 0.157 0.070 0.000 0.788 0.088 0.247 0.096
SCARF 0.696 0.138 0.180 0.004 0.810 0.084 0.291 0.204
SUPPORT SEER
Method CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$ CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$
NI 0.612 0.204 0.149 0.000 0.842 0.014 0.975 0.999
SCARF 0.625 0.195 0.155 0.000 0.854 0.009 0.960 0.999
@@ Appendix C Additional Results
Table 5: Comparison of time-dependent C-index.
METABRIC NWTCO
Methods 25% 50% 75% 25% 50% 75%
CoxPH $0.653_{\pm 0.024}$ $0.645_{\pm 0.018}$ $0.637_{\pm 0.018}$ $0.718_{\pm 0.026}$ $0.711_{\pm 0.025}$ $0.706_{\pm 0.026}$
DeepSurv $0.643_{\pm 0.034}$ $0.620_{\pm 0.024}$ $0.607_{\pm 0.029}$ $0.640_{\pm 0.081}$ $0.636_{\pm 0.079}$ $0.634_{\pm 0.078}$
DeepHit $0.605_{\pm 0.045}$ $0.659_{\pm 0.044}$ $0.603_{\pm 0.032}$ $0.717_{\pm 0.023}$ $\textbf{0.723}_{\pm\textbf{0.023}}$ $0.704_{\pm 0.024}$
DRSA $0.639_{\pm 0.030}$ $0.605_{\pm 0.020}$ $0.592_{\pm 0.022}$ $0.713_{\pm 0.020}$ $0.706_{\pm 0.019}$ $0.704_{\pm 0.019}$
DCS $0.646_{\pm 0.030}$ $0.615_{\pm 0.023}$ $0.606_{\pm 0.024}$ $0.675_{\pm 0.026}$ $0.657_{\pm 0.020}$ $0.643_{\pm 0.029}$
X-CAL $0.661_{\pm 0.025}$ $0.632_{\pm 0.018}$ $0.619_{\pm 0.019}$ $0.687_{\pm 0.026}$ $0.650_{\pm 0.029}$ $0.628_{\pm 0.034}$
ConSurv $\textbf{0.686}_{\pm\textbf{0.028}}$ $\textbf{0.660}_{\pm\textbf{0.021}}$ $\textbf{0.645}_{\pm\textbf{0.023}}$ $\textbf{0.718}_{\pm\textbf{0.025}}$ $0.709_{\pm 0.024}$ $\textbf{0.705}_{\pm\textbf{0.025}}$
GBSG FLCHAIN
Methods 25% 50% 75% 25% 50% 75%
CoxPH $0.714_{\pm 0.027}$ $0.678_{\pm 0.019}$ $0.664_{\pm 0.018}$ $0.796_{\pm 0.019}$ $\textbf{0.793}_{\pm\textbf{0.015}}$ $0.790_{\pm 0.014}$
DeepSurv $0.700_{\pm 0.056}$ $0.668_{\pm 0.047}$ $0.654_{\pm 0.043}$ $0.787_{\pm 0.028}$ $0.783_{\pm 0.027}$ $0.780_{\pm 0.028}$
DeepHit $0.630_{\pm 0.045}$ $0.626_{\pm 0.030}$ $0.609_{\pm 0.028}$ $0.796_{\pm 0.020}$ $\textbf{0.793}_{\pm\textbf{0.014}}$ $0.785_{\pm 0.014}$
DRSA $0.725_{\pm 0.687}$ $0.687_{\pm 0.015}$ $0.671_{\pm 0.016}$ $0.776_{\pm 0.023}$ $0.776_{\pm 0.016}$ $0.772_{\pm 0.013}$
DCS $0.720_{\pm 0.025}$ $0.687_{\pm 0.015}$ $0.674_{\pm 0.017}$ $0.788_{\pm 0.024}$ $0.783_{\pm 0.019}$ $0.779_{\pm 0.018}$
X-CAL $\textbf{0.727}_{\pm\textbf{0.025}}$ $0.680_{\pm 0.016}$ $0.676_{\pm 0.017}$ $0.793_{\pm 0.021}$ $0.788_{\pm 0.015}$ $0.784_{\pm 0.016}$
ConSurv $0.723_{\pm 0.030}$ $\textbf{0.689}_{\pm\textbf{0.019}}$ $\textbf{0.676}_{\pm\textbf{0.020}}$ $\textbf{0.796}_{\pm\textbf{0.022}}$ $0.791_{\pm 0.017}$ $\textbf{0.786}_{\pm\textbf{0.015}}$
SUPPORT SEER
Methods 25% 50% 75% 25% 50% 75%
CoxPH $0.603_{\pm 0.006}$ $0.607_{\pm 0.006}$ $0.609_{\pm 0.006}$ $0.865_{\pm 0.018}$ $0.890_{\pm 0.023}$ $0.855_{\pm 0.022}$
DeepSurv $0.603_{\pm 0.009}$ $0.605_{\pm 0.009}$ $0.608_{\pm 0.010}$ $0.838_{\pm 0.007}$ $0.835_{\pm 0.021}$ $0.836_{\pm 0.002}$
DeepHit $0.516_{\pm 0.007}$ $0.489_{\pm 0.008}$ $0.521_{\pm 0.007}$ $0.847_{\pm 0.016}$ $0.841_{\pm 0.020}$ $0.827_{\pm 0.020}$
DRSA $0.586_{\pm 0.009}$ $0.583_{\pm 0.007}$ $0.580_{\pm 0.009}$ $0.840_{\pm 0.073}$ $0.839_{\pm 0.077}$ $0.820_{\pm 0.083}$
DCS $0.604_{\pm 0.008}$ $0.602_{\pm 0.009}$ $0.602_{\pm 0.009}$ $0.864_{\pm 0.019}$ $0.863_{\pm 0.023}$ $0.858_{\pm 0.023}$
X-CAL $0.610_{\pm 0.006}$ $0.608_{\pm 0.006}$ $0.608_{\pm 0.007}$ $0.840_{\pm 0.023}$ $0.839_{\pm 0.019}$ $0.833_{\pm 0.018}$
ConSurv $\textbf{0.616}_{\pm\textbf{0.006}}$ $\textbf{0.617}_{\pm\textbf{0.006}}$ $\textbf{0.618}_{\pm\textbf{0.006}}$ $\textbf{0.870}_{\pm\textbf{0.017}}$ $\textbf{0.866}_{\pm\textbf{0.020}}$ $\textbf{0.862}_{\pm\textbf{0.020}}$
@@ C.1 Performance on Time-dependent C-index and Brier score
C.2 Effect of Contrastive Learning
Figure 5: t-SNE visualization for latent representations learned with $\mathcal{L}_{\text{{NLL}}}$ only, $\mathcal{L}_{\text{{SNCE}}}$ only, and ConSurv for the FLCHAIN and SUPPORT datasets, colored by event times (for uncensored samples).
@@ C.2 Effect of Contrastive Learning
METABRIC NWTCO
Methods 25% 50% 75% 25% 50% 75%
CoxPH $0.167_{\pm 0.009}$ $\textbf{0.215}_{\pm\textbf{0.014}}$ $\textbf{0.236}_{\pm\textbf{0.038}}$ $\textbf{0.111}_{\pm\textbf{0.009}}$ $0.114_{\pm 0.008}$ $0.114_{\pm 0.008}$
DeepSurv $0.173_{\pm 0.010}$ $0.228_{\pm 0.016}$ $0.249_{\pm 0.048}$ $0.121_{\pm 0.011}$ $0.124_{\pm 0.011}$ $0.125_{\pm 0.012}$
DeepHit $0.181_{\pm 0.009}$ $0.256_{\pm 0.010}$ $0.263_{\pm 0.026}$ $0.151_{\pm 0.009}$ $0.143_{\pm 0.008}$ $0.154_{\pm 0.009}$
DRSA $0.264_{\pm 0.041}$ $0.343_{\pm 0.038}$ $0.337_{\pm 0.079}$ $0.281_{\pm 0.030}$ $0.294_{\pm 0.039}$ $0.310_{\pm 0.051}$
DCS $0.190_{\pm 0.020}$ $0.265_{\pm 0.038}$ $0.293_{\pm 0.083}$ $0.119_{\pm 0.010}$ $0.107_{\pm 0.011}$ $0.122_{\pm 0.024}$
X-CAL $0.180_{\pm 0.012}$ $0.234_{\pm 0.021}$ $0.254_{\pm 0.057}$ $0.113_{\pm 0.009}$ $\textbf{0.106}_{\pm\textbf{0.013}}$ $0.137_{\pm 0.036}$
ConSurv $\textbf{0.166}_{\pm\textbf{0.009}}$ $0.221_{\pm 0.015}$ $0.255_{\pm 0.045}$ $0.111_{\pm 0.009}$ $0.113_{\pm 0.008}$ $\textbf{0.113}_{\pm\textbf{0.009}}$
GBSG FLCHAIN
Methods 25% 50% 75% 25% 50% 75%
CoxPH $\textbf{0.145}_{\pm\textbf{0.010}}$ $0.211_{\pm 0.007}$ $0.219_{\pm 0.010}$ $\textbf{0.074}_{\pm\textbf{0.006}}$ $\textbf{0.112}_{\pm\textbf{0.006}}$ $\textbf{0.133}_{\pm\textbf{0.007}}$
DeepSurv $0.146_{\pm 0.011}$ $0.212_{\pm 0.009}$ $0.221_{\pm 0.013}$ $0.076_{\pm 0.007}$ $0.114_{\pm 0.009}$ $0.137_{\pm 0.011}$
DeepHit $0.168_{\pm 0.012}$ $0.242_{\pm 0.008}$ $0.250_{\pm 0.006}$ $0.096_{\pm 0.005}$ $0.181_{\pm 0.004}$ $0.244_{\pm 0.004}$
DRSA $0.210_{\pm 0.016}$ $0.322_{\pm 0.022}$ $0.349_{\pm 0.020}$ $0.141_{\pm 0.013}$ $0.251_{\pm 0.027}$ $0.323_{\pm 0.031}$
DCS $0.153_{\pm 0.010}$ $0.210_{\pm 0.006}$ $0.218_{\pm 0.010}$ $0.079_{\pm 0.008}$ $0.119_{\pm 0.008}$ $0.147_{\pm 0.014}$
X-CAL $0.150_{\pm 0.011}$ $\textbf{0.207}_{\pm\textbf{0.007}}$ $\textbf{0.217}_{\pm\textbf{0.012}}$ $0.080_{\pm 0.005}$ $0.117_{\pm 0.007}$ $0.144_{\pm 0.012}$
ConSurv $0.146_{\pm 0.009}$ $0.209_{\pm 0.008}$ $\textbf{0.217}_{\pm\textbf{0.010}}$ $0.077_{\pm 0.006}$ $0.115_{\pm 0.007}$ $0.136_{\pm 0.008}$
SUPPORT SEER
Methods 25% 50% 75% 25% 50% 75%
CoxPH $\textbf{0.219}_{\pm\textbf{0.003}}$ $0.197_{\pm 0.008}$ $0.180_{\pm 0.009}$ $\textbf{0.005}_{\pm\textbf{0.001}}$ $\textbf{0.009}_{\pm\textbf{0.001}}$ $\textbf{0.014}_{\pm\textbf{0.001}}$
DeepSurv $0.220_{\pm 0.004}$ $0.198_{\pm 0.009}$ $0.180_{\pm 0.010}$ $\textbf{0.005}_{\pm\textbf{0.001}}$ $0.010_{\pm 0.001}$ $0.016_{\pm 0.001}$
DeepHit $0.340_{\pm 0.005}$ $0.282_{\pm 0.003}$ $0.240_{\pm 0.005}$ $0.016_{\pm 0.000}$ $0.020_{\pm 0.001}$ $0.026_{\pm 0.001}$
DRSA $0.285_{\pm 0.013}$ $0.269_{\pm 0.022}$ $0.249_{\pm 0.021}$ $0.011_{\pm 0.006}$ $0.021_{\pm 0.016}$ $0.034_{\pm 0.026}$
DCS $0.222_{\pm 0.006}$ $0.209_{\pm 0.015}$ $0.202_{\pm 0.017}$ $\textbf{0.005}_{\pm\textbf{0.001}}$ $0.010_{\pm 0.001}$ $0.015_{\pm 0.002}$
X-CAL $0.220_{\pm 0.005}$ $0.207_{\pm 0.016}$ $0.201_{\pm 0.018}$ $0.006_{\pm 0.001}$ $0.011_{\pm 0.002}$ $0.019_{\pm 0.003}$
ConSurv $\textbf{0.219}_{\pm\textbf{0.005}}$ $\textbf{0.195}_{\pm\textbf{0.010}}$ $\textbf{0.177}_{\pm\textbf{0.010}}$ $\textbf{0.005}_{\pm\textbf{0.001}}$ $\textbf{0.009}_{\pm\textbf{0.001}}$ $0.015_{\pm 0.001}$
C.3 Sensitive Analysis